OpenStudy (anonymous):

A piece of iron weighing 62.3g was heated to 129*C and then dropped into a container containing 209.4g of water at 23.1*C. What is the final temp of the water when thermal equilibrium is reached? C(H2O)=4.184J/gK C(Fe)=0.449J/gK

OpenStudy (anonymous):

I know I am using q=cm (delta)T and that q(Fe)+q(H20)=0

OpenStudy (anonymous):

Just not sure where to plug everything in.

OpenStudy (anonymous):

I keep getting 3-6*C off from the answer which I know is 26.4*C or 299.6K

OpenStudy (aaronq):

equate them to each other, deltaT= Tf- Ti iron water mC(Tf- Ti) = mC(Tf- Ti)

OpenStudy (anonymous):

So I would set it up like (0.449)(62.4)(x-129.2)=(4.184)(209.4)(23.1)

OpenStudy (anonymous):

I am getting 19.6*C

OpenStudy (anonymous):

My teacher mentioned in her lecture that qFe=-qH2O

OpenStudy (anonymous):

And that qFe=-CH20m(tf-ti) But I am not getting the right answer it that either.

OpenStudy (aaronq):

sorry i did forget the - sign (0.449)(62.4)(x-129.2)=- (4.184)(209.4)(x-23.1)

OpenStudy (aaronq):

that was close, but you didn't add the x on the water side

OpenStudy (anonymous):

Oh ok. So C for water is negative but C for iron stays positive. doing it that way gets me the right answer. Thank you so much. This is a study problem for a test we have on Monday and I just couldn't figure it out. Thanks again.

OpenStudy (aaronq):

no problem. well it's not the C that is negative it's because the heat, q, given off by the iron is fully absorbed by the water so yeah \[q _{Fe}=-q _{H _{2}0}\]